what is the value of x if the slope of the line through A(3,-2) B(x,-5)is -3?
Answer by Cromlix(4381)

Using y2 - y1/x2 - x1 = -3
-5 -(-2)/x - 3 = -3
-5 + 2/ x - 3 = -3
-3/ x - 3 = -3
Cross multiply
-3 = -3(x - 3)
-3 = -3x + 9
-3x = -3 - 9
-3x = -12
x = 4
Proof:
y2 - y1/x2 - x1 = -3
-5 -(-2)/4 - 3 = -3
-3/1 = -3.
Hope this helps.
:-)